# Binding Affinity Prediction: From Conventional to Machine Learning-Based Approaches

### v1

@@ Abstract
Protein-ligand binding is the process by which a small molecule (drug or inhibitor) attaches to a target protein. The binding affinity, which refers to the strength of this interaction, is central to many important problems in bioinformatics such as drug design. An extensive amount of work has been devoted to predicting binding affinity over the past decades due to its significance. In this paper, we review all significant recent works, focusing on the methods, features, and benchmark datasets. We have observed a rising trend in the use of traditional machine learning and deep learning models for predicting binding affinity, accompanied by an increasing amount of data on proteins and small drug-like molecules. While prediction results are constantly improving, we also identify several open questions and potential directions that remain unexplored in the field. This paper could serve as an excellent starting point for machine learning researchers who wish to engage in the study of binding affinity, or for anyone with general interests in machine learning, drug discovery, and bioinformatics.
@@ 1 Introduction
Protein-ligand binding (Clyde et al., 2023) refers to the process as shown in Fig. 1 by which ligands—usually small molecules, ions, or proteins—generate signals by binding to the active sites of target proteins through intermolecular forces. This binding typically changes the conformation of target proteins, which then results in the realization, modulation, or alteration of protein functions. Therefore, protein-ligand binding plays a central role in most, if not all, important life processes. For example, oxygen molecules are bound and carried through the human body by proteins like hemoglobin, and then utilized for energy production, while nonsteroidal anti-inflammatory drugs (NSAIDs) like ibuprofen work by inhibiting the functionality of the cyclooxygenase (COX) enzyme that thus reducing the release of pain-causing substances in the body.
@@
The concept and importance of binding affinity prediction were first addressed in Böhm (1994): given the 3D structures of a target protein and a potential ligand, the objective is to predict the binding constant of such a complex, along with the most probable binding pose candidates. The prediction of the binding site (the set of protein residues that have at least one non-hydrogen atom within 4.0 Å of a ligand’s non-hydrogen atom (Khazanov and Carlson, 2013)) and affinity (binding constants such as inhibition or dissociation constants, or the concentration at 50% inhibition) are usually divided into two separate but related stages (Ballester and Mitchell, 2010a).
Figure 1: Binding Example in Surface View with Protein PDB: 10GS
@@
One notable motivation for constructing a good binding affinity predictor (or scoring function, as called in some earlier work) is the essential role that it plays in drug discovery (Liu et al., 2023, 2024a) and virtual screening (Meng et al., 2011; Pinzi and Rastelli, 2019; Sadybekov and Katritch, 2023). Traditional drug discovery essentially involves a process of trial and error. However, with a functional binding affinity predictor, we can significantly reduce the number of experimental trials by focusing only on drugs estimated to be more effective, thus potentially saving large amounts of both time and money. This motivation is becoming increasingly relevant with the rising costs of drug discovery (Scannell et al., 2012; Hay et al., 2014; Tiwari et al., 2023) and the declining rate of drug approvals over recent decades (Dickson and Gagnon, 2004; Paul et al., 2010). Beyond its economic significance in drug development, binding affinity prediction also provides tools, methods, and insights into many other areas of research in biochemistry, pharmaceutical research, and scientific computation.
Despite the increasing motivation and extensive amount of work, binding affinity prediction remains challenging. First and foremost, our understanding of chemistry is not nearly as sufficient as it needs to be, which often leads to sub-optimal human-engineered features for binding affinity prediction, as shown in Ballester and Mitchell (2010a) and Ballester et al. (2014). Furthermore, synthetic datasets remain undesirable for learning, because computationally generated complexes are either inaccurate or too costly to generate in terms of time and computational resources.
@@
Currently, the datasets used, which are usually obtained experimentally for the aforementioned reason, are each limited in various ways: (1) the total number of data points is increasing but still far from sufficient for large-scale data mining due to the vast time and money required to study each complex; (2) the accuracy of experimental data may be limited by the measurement methods used (Su et al., 2018; Pantsar and Poso, 2018); and (3) samples are often biased toward complexes that have the correct poses and good binding constants, meaning that protein-ligand pairs that fail to bind or have relatively low binding constants are not readily obtainable and therefore are unavailable for learning. All these limitations make the search for a good protein-ligand binding affinity predictor extremely challenging.
@@
To make things more complicated, there are different aspects/sub-problems associated with the prediction of binding constants. Li et al. (2014) and Liu et al. (2017) note four different types of predictions related to binding affinity: (1) scoring, which predicts the binding constant; (2) ranking, which ranks different potential ligands of a target; (3) docking, which predicts the best binding pose from multiple decoys; and (4) screening, which predicts the best ligand from multiple decoys. These sub-problems are closely related to each other, to the extent that solving one without addressing the others adds to the existing challenges. Thus, despite all the work and effort put into protein-ligand binding affinity prediction, the results remain sub-optimal. It is often the case that a predictor excels in scoring but fails miserably in docking or screening, rendering the whole system almost useless for virtual screening, one of the ultimate goals of binding affinity prediction.
Figure 2: Binding affinity pipeline (in cartoon view): Computational binding affinity measurements are typically performed using molecular docking simulations. The simulations search for an optimal binding pose for each compound and produces a score for each compound. These scores are then used to rank compounds with respect to each other.
Nevertheless, all previous works can be roughly divided into three distinct categories based on the methods used for binding affinity prediction: conventional, traditional machine-learning-based, and deep-learning-based. Conventional methods, as shown in Fig. 2, are typically based purely on knowledge and assumptions about chemistry and quantum mechanics, often formulated as a set of rules or equations that combine different factors into a single binding affinity result. These methods are usually rigid and tend to work well only in specific scenarios, such as with a single protein family (Bender et al., 2021).
@@
Since 2005, many traditional machine learning methods (such as random forest and support vector machines) have been applied to human-engineered features extracted from complex structures, achieving some improvement over conventional approaches. Machine learning methods have proven to be less rigid and often more accurate, especially in terms of binding affinity scoring and ranking (Ballester and Mitchell, 2010b; Zilian and Sotriffer, 2013; Li et al., 2011) A decade later, as the number of protein-ligand samples in standard benchmarks like PDBbind (Wang et al., 2004a), MOAD (Hu et al., 2005), and PDB (Burley et al., 2017) increased, deep learning began to dominate. This approach typically relies less on human-engineered features, if at all, and its learning potential could greatly increase with the continued accumulation of data.
@@
This paper covers important work on protein-ligand binding affinity prediction, from the early 2000s to the present, along with commonly used datasets and evaluation metrics. The remainder of the paper is organized as follows: Section 2 introduces important datasets and benchmarks, including data specifications and various datasets. Section 3 discusses methodology, in two categories: conventional and machine learning-based methods. In Section 4, we introduce various evaluation approaches, including scoring power, docking power, and ranking power. Finally, we conclude with a discussion of open questions and possible directions that remain unexplored in the field.
@@ 2 Datasets and Benchmarks
There are many datasets and benchmarks used for the study of protein-ligand binding affinity prediction, each focusing on different aspects of the problem. Here, we first discuss in Section 2.1 the specifications for binding affinity data, such as resolution, measurement, binding constant, and concentration. These factors are essential when selecting the right dataset for a specific goal. In Section 2.2, we review some of the most commonly used datasets and benchmarks directly applied in the learning and prediction of protein-ligand binding affinity. These datasets usually have binding constants such as $K_{i}$, $K_{d}$, or $IC_{50}$ associated with most, if not all, protein-ligand complexes. We also note other datasets of protein-ligand binding affinity not directly used for model training and evaluation. Even though these datasets are not directly applicable for training and evaluation, they can be useful for pre-training and may provide additional information, methods, and insights. All discussed datasets are listed in Table 1.
@@ 2.1.1 Structure
The structure of a protein-ligand complex refers to its conformation, or in other words, the spatial arrangement of its atoms. This is an important specification or feature of binding affinity data, as it opens up numerous possibilities for feature engineering and learning methods. For instance, with given structures, we can parameterize the energy terms and atom-pair distances and use them for conventional scoring or learning with methods such as random forests (Ballester and Mitchell, 2010a), support vector machines (Kinnings et al., 2011), and neural networks (Durrant and McCammon, 2010, 2011). On the other hand, measuring structures is rather expensive (more than \$1,000 USD for solving a structure by crystal diffraction) compared to sequence identity, along with the additional costs of data storage and manipulation. Still, the advantages of having a well-measured structure significantly outweigh the disadvantages, and we are observing a rising trend of more protein-ligand complexes with structures being added to PDB (Protein Data Bank) and other databases.
@@ 2.1.2 Resolution
Resolution, in the context of protein-ligand complex structures, refers to the distance corresponding to the smallest observable feature in the measured structure (Guterres and Im, 2020). In other words, two objects within this distance will be perceived as one and thus rendered indistinguishable. Generally, structures with a resolution finer than 1–1.2 Å are considered high-resolution, while those with a resolution lower than 3.0 Å only outline the basic contours of the protein chains. Resolution is an essential measurement for the quality of data and is used as a selection criterion for many binding affinity databases, such as CASF and Binding MOAD.
@@ 2.1.3 Structure Determination Techniques
Cryo-EM uses a mechanism called electron scattering, where electron beams pass through an instantly cooled protein solution, scatter into a lens, and are then converted into a series of 2D images on a detector. These images are processed by reconstruction software to obtain the 3D structure. Cryo-EM is becoming increasingly popular because it works well with larger proteins, has an easier sample preparation process, and yields 3D structures that are much closer to their native state than those obtained from X-ray crystallography. However, the resolution of the structures can be relatively low due to the high level of noise and unknown orientations.
@@ 2.1.4 Binding Affinity Measurement
Roughly speaking, binding affinity measures how “tightly” a ligand binds to a target protein, which is determined with radioligand-binding experiments (Haylett, 2003). A radioligand-binding experiment can produce two kinds of binding affinity data: (1) binding constants such as dissociation constant ($K_{d}$) and inhibition constant ($K_{i}$), which are usually interchangeable in the context of protein-ligand binding; and (2) concentration terms like the half maximal inhibitory concentration (${IC}_{50}$) or half maximal effective concentration (${EC}_{50}$). In binding affinity databases, each protein-ligand complex is typically associated with one or more binding data of $K_{d}$, $K_{i}$, and ${IC}_{50}$, with rare exception. In certain databases and studies, like the “refined set” in PDBbind, complexes with concentration terms only will be filtered out for better data quality. This is due to the fact that ${IC}_{50}$ or ${EC}_{50}$ values depend on the radioligand concentration and can vary between different experiments, unlike constant terms. If necessary or useful, the inhibitory constant $K_{i}$ value can be estimated from $IC_{50}$ via the Cheng-Prusoff equations:
@@ 2.2 Binding Affinity Datasets
PDBbind. Starting in 2004, this dataset has been updated annually (Wang et al., 2005; Liu et al., 2015, 2017). It collects not only protein-ligand complexes but also protein-protein, protein-nucleic acid, and nucleic acid-ligand complexes from the Protein Data Bank (PDB) (Berman et al., 2000), without any restrictions on resolution, binding data, or structural measurement techniques. An important feature of PDBbind is the refined set, selected based on (1) the quality of the complex structure, (2) the quality of binding data, and (3) the nature of the complex (molecular weight, atom types, surface area, etc.). See the supporting material in (Liu et al., 2017) for more details on the selection criteria. However, according to the authors, the refined set, despite its much higher quality compared to the general set, should not be regarded as a high-quality dataset but rather as a collection of complex samples that lack any obvious problems. With the increasing number of data points, subsets with quality control, and easy accessibility, PDBbind is becoming one of the most used datasets in the research of protein-ligand binding affinity, if not the most used.
@@
CASF (Comparative Assessment of Scoring Functions). CASF (Cheng et al., 2009; Li et al., 2018; Su et al., 2018) was proposed along with the development of PDBbind. Originally named the core set in PDBbind, CASF was selected from the refined set and serves as the test set/benchmark for all the protein-ligand binding predictor. To ensure that the samples in the test set are diverse but not redundant, all complexes in the refined set are first clustered based on sequence similarity. Then, for each cluster, three complexes, each with high, medium, and low binding affinity, are selected for inclusion in CASF. With this type of data selection, we can not only test regression on binding affinity (scoring) but also ranking (three different complexes in the same cluster). Additionally, with the given decoys of poses and ligands in CASF, we can evaluate the docking and screening power of a predictor. Overall, CASF provides researchers with a comprehensive and easy-to-use benchmark for protein-ligand binding affinity predictors.
@@
Binding MOAD (Mother of All Databases). Binding MOAD (Hu et al., 2005; Ahmed et al., 2015; Smith et al., 2019) is another commonly used dataset for binding studies. Proposed in 2005 and updates annually, Binding MOAD is a subset of PDB, aiming to be the largest collection of high-quality protein-ligand complexes annotated with experimentally determined binding affinity. Despite sharing the same primary data source (PDB) and a similar goal, Binding MOAD has much in common with PDB. The main differences that set these two apart for users are: (1) Binding MOAD sets the data selection criteria somewhere between PDBbind’s general set and refined set. Notably, Binding MOAD only contains valid protein-ligand complexes with crystal structures of 2.5 Åresolution or better. Additionally, protein sequences are clustered to avoid redundancy of the data, similar to the refined set in PDBbind. (2) Binding MOAD contains complexes without binding data, which is not the case in PDBbind. In fact, Binding MOAD contains only 12,098 binding data of 32,747 complexes. This means that binding MOAD, while having more complexes with structures than PDBbind, contains fewer samples with binding data.
@@
BindingDB. Launched in 2000 and updated weekly as a web-accessible database, BindingDB (Liu et al., 2007; Gilson et al., 2016) collects protein-compound pairs with associated affinity data primarily from scientific articles and, increasingly, patents. It also accepts direct deposition of binding data by users. Newly curated data from these original sources are checked by BindingDB staff to ensure reliability. Additionally, it integrates complexes with affinity measurements from PubChem (Wang et al., 2009), ChEMBL (Bento et al., 2014), PDSP Ki (Roth et al., 2000), and CSAR (Carlson and Dunbar Jr, 2011), allowing users to access all the data through the unified interface of BindingDB (Gilson et al., 2016). There are two important differences between BindingDB and other large binding affinity datasets such as PDBbind and Binding MOAD. First, although most binding data are experimentally determined, BindingDB contains some complexes that are computationally generated with clear labels. Secondly and more importantly, only a fraction of BindingDB’s affinity data has associated structures. In fact, according to their website, there are 2,291 protein-ligand crystal structures available in BindingDB with 100% sequence identity and 5,816 crystal structures with protein sequence identity as low as 85%. The variety of data sources and the aforementioned differences between BindingDB and other databases make it the largest protein-ligand binding affinity database in terms of the number of complexes.
@@
DUD-E (A Directory of Useful Decoys: Enhanced). To combat data bias and benchmark ligand enrichment against challenging decoys, DUD (Huang et al., 2006) (Directory of Useful Decoys) and its successor DUD-E (Mysinger et al., 2012) (A Directory of Useful Decoys: Enhanced) were proposed in 2006 and 2012, respectively. The dataset was constructed by sourcing ligands from ChEMBL, structural data from PDB, and potential decoys from ZINC. The docking poses were generated with DOCK (Ewing et al., 2001; Allen et al., 2015). Immediately after being proposed and constructed, DUD became the gold standard for the evaluation of virtual screening methods (Réau et al., 2018). However, as pointed out in Good and Oprea (2008), the original DUD dataset had a number of issues, such as target selection and analogue bias. To address these issues, the authors of DUD released DUD-E, which added net charge into consideration during decoy selection. In DUD-E, each of the 102 target proteins has 50 decoys that are similar to the ligand in terms of 1D physico-chemical properties, such as molecular weight and calculated LogP, to remove dataset bias, but differ in 2D topology to be likely non-binders. Nevertheless, DUD-E remains unsatisfactory, as shown in multiple works (Xia et al., 2015; Chen et al., 2019; Sieg et al., 2019), indicating that hidden bias in the dataset might still exist, leading to misleadingly good benchmarking results for many docking methods. A good experiment to reveal this bias is to featurize the decoy ligands and plot them using 2D or 3D t-SNE to see if the decoy and real ligands separate.
@@
BioLiP (Yang et al., 2013a) and its successor BioLiP2 (Zhang et al., 2024) (collectively referred to as BioLiP) are semi-manually curated databases for protein-ligand interactions. Founded in 2013 and updated weekly, BioLiP is designed to focus only on biologically relevant ligands, meaning they are not simply additives for protein purification and/or crystallization. This is achieved through automatic procedures and manual verification. In addition to annotations of binding affinity, the dataset also contains relevant functional annotations of the complexes. It sources structures from PDB and binding data from multiple databases such as Binding MOAD, PDBbind, BindingDB, and literature. It is worth noting that the main focus of BioLiP is binding site prediction, similar to FireDB (Maietta et al., 2014) and LigASite (Dessailly et al., 2007), and binding affinity is one of the optional fields for each entry in the database. Only about 5% of the protein-ligand complexes have associated binding affinity.

### v2

@@ Abstract
Protein-ligand binding is the process by which a small molecule (drug or inhibitor) attaches to a target protein. Binding affinity, which characterizes the strength of biomolecular interactions, is essential for tackling diverse challenges in life sciences, including therapeutic design, protein engineering, enzyme optimization, and elucidating biological mechanisms. Much work has been devoted to predicting binding affinity over the past decades. Here, we review recent significant works, with a focus on methods, evaluation strategies, and benchmark datasets. We note growing use of both traditional machine learning and deep learning models for predicting binding affinity, accompanied by an increasing amount of data on proteins and small drug-like molecules. With improved predictive performance and the FDA’s phasing out of animal testing, AI-driven in silico models, such as AI virtual cells (AIVCs), are poised to advance binding affinity prediction; reciprocally, progress in building binding affinity predictors can refine AIVCs. Future efforts in binding affinity prediction and AI-driven in silico models can enhance the simulation of temporal dynamics, cell-type specificity, and multi-omics integration to support more accurate and personalized outcomes.
@@ 1 Introduction
Protein-ligand binding (Clyde et al., 2023) refers to the process by which ligands—usually small molecules, ions, or proteins—generate signals by binding to the active sites of target proteins through intermolecular forces (Fig. 1). This process typically changes the conformation of target proteins, which then results in the realization, modulation, or alteration of protein functions. As such, protein–ligand binding affinity is a fundamental parameter in understanding and optimizing molecular interactions, with broad applications across biomedical research and drug development. In small-molecule drug discovery, it guides hit identification, lead optimization, and candidate selection to ensure strong, selective binding to target proteins. In biologics and therapeutic protein design, it informs the engineering of enzymes, receptors, and antibodies for improved specificity and potency. In diagnostics, it underlies the development of high-sensitivity biosensors and molecular probes for detecting disease biomarkers. In precision medicine, protein–ligand affinity data help predict patient-specific drug responses, uncover mechanisms of resistance, and tailor treatments, while also finding utility in agricultural and environmental applications through the design of selective binding agents.
@@
Binding affinity prediction was introduced in Böhm (1994): Given the 3D structures of a target protein and a potential ligand, predict the binding constant of the complex along with the most probable binding pose candidates. The prediction of the binding site (i.e., the set of protein residues that have at least one non-hydrogen atom within 4.0 Å of a ligand’s non-hydrogen atom (Khazanov and Carlson, 2013)) and affinity (i.e., binding constants such as inhibition or dissociation constants, or the concentration at 50% inhibition) are usually divided into two separate but related stages (Ballester and Mitchell, 2010a).
Figure 1: Overview of binding affinity. (A) Binding affinity pipeline. Computational binding affinity measurements are typically performed using molecular docking simulations as a surrogate. For each compound, the simulations search for an optimal binding pose and produces a score. These scores are then used to rank compounds with respect to each other.(B) Binding Example in Surface View with Protein PDB: 10GS. (C) Detailed pipline. For a given protein and a set of candidate molecules, we perform both single-site docking and multi-site docking. The top row illustrates the search for optimal molecules and corresponding binding poses across multiple potential binding pockets, while the bottom row depicts the docking of various molecules into a single predefined pocket. By combining these two strategies, we aim to identify the most favorable molecule, binding pose, and docking score. (D) Application domains. Binding affinity is crucial across drug discovery, biologics design, diagnostics, and precision medicine. It guides the identification and optimization of molecules—such as small drugs, antibodies, or probes—for strong and selective target binding.
@@
One notable motivation for constructing a good binding affinity predictor (or scoring function) is the essential role that it plays in drug discovery (Liu et al., 2023, 2024b) and virtual screening (Meng et al., 2011; Pinzi and Rastelli, 2019; Sadybekov and Katritch, 2023). Traditional drug discovery involves a process of trial and error; however, with a functional binding affinity predictor, we can significantly reduce the number of experimental trials by focusing only on drugs estimated to be more effective, thereby reducing costs (Scannell et al., 2012; Hay et al., 2014; Tiwari et al., 2023). Beyond its economic significance in drug development, binding affinity prediction also provides tools, methods, and insights into many other areas of research in biochemistry, pharmaceutical research, and scientific computation.
@@
Despite the increasing motivation and extensive amount of work, binding affinity prediction remains challenging. First, our understanding of chemistry is not comprehensive, which often leads to sub-optimal human-engineered features for binding affinity prediction (Ballester and Mitchell, 2010a; Ballester et al., 2014). Secondly, synthetic datasets remain undesirable for learning because computationally generated complexes are either inaccurate or too costly to generate in terms of time and computational resources. The data sets currently being used, which are usually obtained experimentally, are limited in various ways: (1) the total number of data points is increasing but still far from sufficient for large-scale data mining due to the vast time and money required to study each complex; (2) the precision of the experimental data may be limited by the measurement methods used (Su et al., 2018; Pantsar and Poso, 2018); and (3) samples are often biased toward complexes that have the correct poses and good binding constants, meaning that protein-ligand pairs that do not bind or have relatively low binding constants are not readily obtainable. These limitations make the search for a good protein-ligand binding affinity predictor extremely challenging.
@@
Further, there are different aspects/sub-problems associated with the prediction of binding constants. Li et al. (2014a) and Liu et al. (2017) note four different types of predictions related to binding affinity: (1) scoring, which predicts the binding constant; (2) rank ordering, which ranks different potential ligands of a target; (3) docking, which predicts the best binding pose from multiple decoys; and (4) screening, which predicts the best ligand from multiple decoys. These sub-problems are closely related to each other, to the extent that solving one without addressing the others adds to the existing challenges. Thus, despite existing efforts to improve protein-ligand binding affinity prediction, the results remain sub-optimal. It is often the case that a predictor excels in scoring for known targets but fails miserably in docking or screening for new targets, rendering the whole system almost useless for virtual screening, one of the ultimate goals of binding affinity prediction.
@@
Existing methods for binding affinity prediction can be roughly divided into three distinct categories: conventional, traditional machine-learning-based, and deep-learning-based. Conventional methods are typically based on ab initio quantum mechanical calculations or empirical approaches derived from experimental data, often formulated as physics-based models or parametric equations that predict binding affinity(Fig. 1(A)). These methods are usually rigid and tend to work well only in specific scenarios, such as with a single protein family (Bender et al., 2021). Since 2005, many traditional machine learning methods have been applied to human-engineered features extracted from complex structures, achieving some improvement over conventional approaches. Machine learning methods have proven to be less rigid and often more accurate, especially in terms of binding affinity scoring and ranking (Ballester and Mitchell, 2010b; Zilian and Sotriffer, 2013; Li et al., 2011) A decade later, as the number of protein-ligand samples in standard benchmarks like PDBbind (Wang et al., 2004a), MOAD (Hu et al., 2005), and PDB (Burley et al., 2017) increased, deep learning began to dominate. This approach typically relies less on human-engineered features, if at all, and its learning potential could greatly increase with the continued accumulation of data.
Looking ahead, the increasing emphasis on in silico drug discovery—accelerated by the FDA’s move to phase out animal testing—positions AI-driven computational models as a transformative force in binding affinity prediction. In particular, emerging AI virtual cells (AIVCs) offer a systems-level framework for modeling molecular interactions in dynamic, cell-specific, and multi-omic contexts. Advances in binding affinity predictors will not only strengthen the molecular foundations of AIVCs but also benefit from the broader simulation capacities that AIVCs provide. This reciprocal relationship underscores a future in which binding affinity prediction and AI-driven in silico modeling co-evolve to enable more accurate, mechanistic, and personalized biomedical insights.
@@
This Review covers foundational research on protein-ligand binding affinity prediction, from the early 2000s to the present, along with commonly used datasets and evaluation metrics. It is organized as follows. Section 2 introduces important datasets and benchmarks, including data specifications and various datasets. Section 4 discusses conventional and machine learning-based methods. Section 3 introduces various evaluation approaches, including scoring power, docking power, and ranking power. Finally, we discuss open questions and research directions that remain unexplored in the field.
@@ 2 Datasets and Benchmarks
There are many datasets and benchmarks used for the study of protein-ligand binding affinity prediction, each focusing on different aspects of the problem. In Section 2.1, we discuss the specifications for binding affinity data, such as resolution, measurement, binding constant, and concentration. These factors are essential when selecting the right dataset for a specific goal. In Section 2.2, we review some of the most commonly used datasets and benchmarks for developing models to predict protein-ligand binding affinity (Table 1). These datasets usually have binding constants such as $K_{i}$, $K_{d}$, or $IC_{50}$ associated with most, if not all, protein-ligand complexes. We also note datasets of protein-ligand binding affinity not directly used for model training and evaluation. While these datasets are not directly applicable for training and evaluation, they can be useful for pre-training and may provide additional information, methods, and insights.
@@ 2.1.1 Structure
The structure of a protein-ligand complex refers to its conformation (or the spatial arrangement of its atoms). This is an important specification or feature of binding affinity data, as it opens up numerous possibilities for feature engineering and learning methods. For instance, with given structures, we can parameterize the energy terms and atom-pair distances and use them for conventional scoring or learning with methods such as random forests (Ballester and Mitchell, 2010a), support vector machines (Kinnings et al., 2011), and neural networks (Durrant and McCammon, 2010, 2011). However, measuring structures is rather expensive (more than \$1,000 USD for solving a structure by crystal x-ray diffraction) compared to sequence identity, along with the additional costs of data storage and manipulation. Still, the advantages of having a well-measured structure significantly outweigh the disadvantages, and we are observing a rising trend of more protein-ligand complexes with structures being added to the Protein Data Bank (PDB) and other databases.
@@ 2.1.2 Resolution
Resolution, in the context of protein-ligand complex structures, refers to the distance corresponding to the smallest observable feature in the measured structure (Guterres and Im, 2020). In other words, two objects within this distance will be perceived as one and thus rendered indistinguishable. Generally, structures with a resolution finer than 1–1.2 Å are considered high-resolution, while those with a resolution lower than 3.0 Å only outline the basic contours of the protein chains. Resolution is an essential measurement for the quality of data and is used as a selection criterion for many binding affinity databases (e.g., CASF, Binding MOAD). Importantly, the effective resolution at the binding site is often more critical than the global resolution of the structure, since small errors in ligand–protein contact regions can disproportionately affect the accuracy of binding affinity estimation and downstream modeling.
@@ 2.1.3 Structure determination techniques
Cryo-EM uses a mechanism called electron scattering, where electron beams pass through an instantly cooled protein solution, scatter into a lens, and are converted into a series of 2D images on a detector. These images are processed by reconstruction software to obtain the 3D structure. Cryo-EM is becoming increasingly popular because it works well with larger proteins, has an easier sample preparation process, and yields 3D structures that are much closer to their native state than those obtained from X-ray crystallography. However, the resolution of the structures can be relatively low due to the high level of noise and unknown orientations.
@@ 2.1.4 Binding affinity measurement
Binding affinity measures how “tightly” a ligand binds to a target protein, which is determined with radioligand-binding experiments (Haylett, 2003). A radioligand-binding experiment can produce two kinds of binding affinity data: (1) binding constants, such as dissociation constant ($K_{d}$) and inhibition constant ($K_{i}$), which are usually interchangeable in the context of protein-ligand binding; and (2) concentration terms, like the half maximal inhibitory concentration (${IC}_{50}$) or half maximal effective concentration (${EC}_{50}$). In binding affinity databases, each protein-ligand complex is typically associated with one or more binding data of $K_{d}$, $K_{i}$, and ${IC}_{50}$. In certain databases and studies, like the refined set in PDBbind, complexes with concentration terms only will be filtered out for better data quality. This is due to the fact that ${IC}_{50}$ or ${EC}_{50}$ values depend on the radioligand concentration and can vary between different experiments, unlike constant terms. If necessary or useful, the inhibitory constant $K_{i}$ value can be estimated from $IC_{50}$ via the Cheng-Prusoff equations:
@@ 2.2 Binding Affinity Datasets
PDBbind. Starting in 2004, PDBbind has been updated annually (Wang et al., 2005; Liu et al., 2015, 2017). It collects not only protein-ligand complexes but also protein-protein, protein-nucleic acid, and nucleic acid-ligand complexes from the Protein Data Bank (PDB) (Berman et al., 2000) without any restrictions on resolution, binding data, or structural measurement techniques. An important feature of PDBbind is the refined set, selected based on (1) the quality of the complex structure, (2) the quality of binding data, and (3) the nature of the complex (e.g., molecular weight, atom types, surface area) (Liu et al., 2017). However, the refined set, despite its much higher quality compared to the general set, should not be regarded as a high-quality dataset but rather as a collection of complex samples that lack any obvious problems (Liu et al., 2017). With the increasing number of data points, subsets with quality control, and easy accessibility, PDBbind is becoming one of the most used datasets in the research of protein-ligand binding affinity.
@@
CASF (Comparative Assessment of Scoring Functions). CASF (Cheng et al., 2009; Li et al., 2018; Su et al., 2018) was introduced along with PDBbind. Originally named the core set in PDBbind, CASF was selected from the refined set and serves as the test set or benchmark for protein-ligand binding predictors. To ensure that the samples in the test set are diverse but not redundant, all complexes in the refined set are first clustered based on sequence similarity. Then, for each cluster, three complexes, each with high, medium, and low binding affinity, are selected for inclusion in CASF. With this type of data selection, we can not only test regression on binding affinity (scoring) but also ranking (three different complexes in the same cluster). Additionally, with the given decoys of poses and ligands in CASF, we can evaluate the docking and screening power of a predictor. Overall, CASF provides researchers with a comprehensive and easy-to-use benchmark for protein-ligand binding affinity predictors.
@@
Binding MOAD (Mother of All Databases). Binding MOAD (Hu et al., 2005; Ahmed et al., 2015; Smith et al., 2019) is another commonly used dataset for binding studies. Proposed in 2005 and updated annually, Binding MOAD is a subset of PDB, aiming to be the largest collection of high-quality protein-ligand complexes annotated with experimentally determined binding affinity. Despite sharing the same primary data source (PDB) and a similar goal, Binding MOAD has much in common with PDB. The main differences that set these two apart for users are: (1) Binding MOAD sets the data selection criteria somewhere between PDBbind’s general set and refined set. Notably, Binding MOAD only contains valid protein-ligand complexes with crystal structures of 2.5 Åresolution or better. Additionally, protein sequences are clustered to avoid redundancy of the data, similar to the refined set in PDBbind. (2) Binding MOAD contains complexes without binding data, which is not the case in PDBbind. In fact, Binding MOAD contains only 12,098 binding data of 32,747 complexes. This means that Binding MOAD, while having more complexes with structures than PDBbind, contains fewer samples with binding data.
@@
BindingDB. Launched in 2000 and updated weekly as a web-accessible database, BindingDB (Liu et al., 2007; Gilson et al., 2016) collects protein-compound pairs with associated affinity data primarily from scientific articles and, increasingly, patents. It also accepts direct deposition of binding data by users. Newly curated data from these original sources are checked by BindingDB staff to ensure reliability. Additionally, it integrates complexes with affinity measurements from PubChem (Wang et al., 2009), ChEMBL (Bento et al., 2014), PDSP Ki (Roth et al., 2000), and CSAR (Carlson and Dunbar Jr, 2011), allowing users to access the data through the unified interface of BindingDB (Gilson et al., 2016). There are two important differences between BindingDB and other large binding affinity datasets, such as PDBbind and Binding MOAD. First, although most binding data are experimentally determined, BindingDB contains some complexes that are computationally generated with clear labels. Secondly and more importantly, only a fraction of BindingDB’s affinity data has associated structures. In fact, according to their website, there are 2,291 protein-ligand crystal structures available in BindingDB with 100% sequence identity and 5,816 crystal structures with protein sequence identity as low as 85%. As such, BindingDB is a protein-ligand binding affinity database with the largest number of complexes.
@@
DUD-E (A Directory of Useful Decoys: Enhanced). To combat data bias and benchmark ligand enrichment against challenging decoys, DUD (Huang et al., 2006) (Directory of Useful Decoys) and its successor DUD-E (Mysinger et al., 2012) (A Directory of Useful Decoys: Enhanced) were proposed in 2006 and 2012, respectively. The dataset was constructed by sourcing ligands from ChEMBL, structural data from PDB, and potential decoys from ZINC. The docking poses were generated with DOCK (Ewing et al., 2001; Allen et al., 2015). Immediately after being proposed and constructed, DUD became the gold standard for the evaluation of virtual screening methods (Réau et al., 2018). However, as pointed out in Good and Oprea (2008), the original DUD dataset had a number of issues, such as target selection and analogue bias. To address these issues, the authors of DUD released DUD-E, which added net charge into consideration during decoy selection. In DUD-E, each of the 102 target proteins has 50 decoys that are similar to the ligand in terms of 1D physico-chemical properties, such as molecular weight and calculated LogP, to remove dataset bias, but differ in 2D topology to be likely non-binders. Nevertheless, DUD-E has been shown to remain unsatisfactory (Xia et al., 2015; Chen et al., 2019; Sieg et al., 2019), indicating that hidden bias in the dataset might still exist, leading to misleadingly good benchmarking results for many docking methods.
@@
BioLiP. BioLiP (Yang et al., 2013a) and its successor BioLiP2 (Zhang et al., 2024) (collectively referred to as BioLiP) are semi-manually curated databases for protein-ligand interactions. Founded in 2013 and updated weekly, BioLiP is designed to focus only on biologically relevant ligands, meaning they are not simply additives for protein purification and/or crystallization. This is achieved through automatic procedures and manual verification. In addition to annotations of binding affinity, the dataset also contains relevant functional annotations of the complexes. It sources structures from PDB and binding data from multiple databases, such as Binding MOAD, PDBbind, BindingDB, and the literature. It is worth noting that the main focus of BioLiP is binding site prediction, similar to FireDB (Maietta et al., 2014) and LigASite (Dessailly et al., 2007). Binding affinity is one of the optional fields for each entry in the database. As such, only about 5% of the protein-ligand complexes have associated binding affinity.